OpenStudy (anonymous):

Which literal equations are equivalent to F = ma? Choose all answers that are correct. A.M=Fa B.M=F/a C.a=M/F D.a=F/M @iGreen @chefducky @dan815 @Perl @bibby @Werewolfprincess @WhiteFangIII Please help fan and Medal!!!!!

TheSmartOne (thesmartone):



OpenStudy (igreen):

Divide m to both sides, then divide a to both sides.

TheSmartOne (thesmartone):

What I said :D

OpenStudy (igreen):

Dividing \(m\) will give us: \(F = ma\) \(\dfrac{F}{m} = \dfrac{ma}{m}\) \(\dfrac{F}{m} = a\) or \(a = \dfrac{F}{m}\) Now can you do the same thing with \(a\)? @WolvesAllTheWay2002

OpenStudy (anonymous):

I'm confuzed

geerky42 (geerky42):

Not really, "Divide m to both sides, THEN divide a to both sides." lol

TheSmartOne (thesmartone):